A farmer's conveyor belt carries hay bales from the ground to the barn loft which is 12 ft from the ground. The conveyor rises at a 60 degree angle. a.)If the bales are moving at a 10 ft/min, how many minutes does it take for them to make it to the loft? b.) how far from the barn does the conveyor belt begin?
Answer by mananth(16946)

A farmer's conveyor belt carries hay bales from the ground to the barn loft which is 12 ft from the ground. The conveyor rises at a 60 degree angle. a.)If the bales are moving at a 10 ft/min, how many minutes does it take for them to make it to the loft? b.) how far from the barn does the conveyor belt begin?
Length of belt = 12/sin 60 =13.856 feet
bales move at 10 ft/m
13.856/10 = .1.39 minutes to move to loft
distance of belt from barn = 12/tan 60 =6.93 feet